Imagine:

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. In parallelogram ABCD, m∠ABD = 83°, m∠BDA = 34°, and m∠BCD = _______- °.

Imagine:

1 attachment
Imagine:

Wait, should I just add 83+34+?=180?

Imagine:

That ?=BCD?

Imagine:

117+?=180

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine Wait, should I just add 83+34+?=180? \(\color{#0cbb34}{\text{End of Quote}}\) naw u just wanna add 83+34=117 180-117=(blank)=BCD

Imagine:

63.

Hoodmemes:

ye

Imagine:

So, 63=BCD?